Please find the degree in these angles.

I have a scalene triangle.

There are sides A, B, and C.

A = (x+2) degree sign.

C= 120 degrees.

B = (x-12)

please find the degree for x m
thank you

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

I have a scalene triangle.
There are sides A, B, and C.
A = (x+2) degree sign.
C= 120 degrees.
B = (x-12)
----------
A + B + C = 180º
x+2 + (x-12) + 120 = 180
2x - 10 = 60
x = 35
----------
A = 37º
B = 23º

Answer by solver91311(24713)



The sum of the angles of any triangle is 180° So, if angle C is 120°, then the sum of angles A and B must be 180 - 120 = 60°



Solve for

John